# Roman Garipov $^{*\,\dagger}$

### v3

@@ 4.3 Inference Benchmarks with vLLM
We compare AutoJudge decoding against lossless vLLM speculative decoding (Leviathan et al., 2023). To ensure fair comparison, we tune draft window size for AutoJudge and baseline independently (see Appendix G). In each configuration, we report the absolute speed in tokens per second (batch size 1), as well as the relative speedup over speculative decoding. We run 1B/8B model pair on a single A100-SXM4-80GB GPU; 8B/70B on 4 A100-SXM4-80G GPUs in tensor-parallel mode. Finally, the 8B/405B runs on 8 H100-SXM5-80GB GPUs with the 405B model loaded in FP8 precision. Additionally, we consider a setup where 8B/70B model pair runs on a single A100-SXM4-80GB GPU with RAM offloading (see Appendix H). We provide additional configuration details in Appendix E.
@@ 5 Conclusion
One promising direction for future research is to focus on open-ended problems such as instruction following or creative writing. While AutoJudge already demonstrates some speedups on open-ended tasks, it was primarily designed for technical problems such as math and programming. It would be interesting to investigate how it can be extended for longform generations with noisy and informal quality criteria. Another possible direction is to combine AutoJudge with additional speculative decoding algorithms, such as speculative decoding with tree-based drafts (Miao et al., 2023; Chen et al., 2024; Svirschevski et al., 2024) or learned drafting heads (Cai et al., 2024; Li et al., 2025).
@@ Acknowledgements
In Section 3.1, we assume that the generation procedure is deterministic, i.e. that the model performs “greedy inference”. In practice, however, many applications work better with stochastic sampling [Holtzman et al., 2020]. However, this has an obvious caveat for Algorithm 1: if the text generation process is stochastic, a token can be deemed important based not on its actual impact on the model outputs, but on the randomness of the decoding procedure.
To generalize our approach for stochastic generation, we take advantage of the well-known Gumbel-max trick [Gumbel, 1954]. To recap, if we add independent Gumbel-distributed random variables to each predicted logit and take the index of the maximum, the probability that a certain index will be chosen is equal to the softmax of the original logits.
@@
As we described earlier, Algorithm 1 is inherently sequential because it searches not for individual important tokens, but for important token combinations. In principle, it is tempting to consider a simpler algorithm that considers each token replacement in isolation and can run in parallel. However, when considering [target_model_gen_0, draft_token, target_model_gen_1] sequences only, a sufficiently strong target model might recover from even a low-quality token and still produce the correct answer. This results in a failure mode where all tokens are individually unimportant, but when all such tokens are jointly replaced with their draft versions, the model fails to produce the correct answer. In our preliminary experiments, when using Llama-3.1-70B-Instruct Touvron et al. [2023] as the target model and Llama-3.2-1B-Instruct as the draft model, fewer than 1% of the tokens were labeled as important with this simplified algorithm, whereas our main Algorithm 1 found substantially more. One interesting guarantee of Algorithm 1 over its naïve counterpart is that whenever draft and target models produce different (non-equivalent) answers to a given prompt, our algorithm will find at least one important token, whereas the naïve algorithm may find none.
@@ Appendix B Additional Details on Classifier Training
As we discussed earlier in Section 3.2, there are several important design choices that can affect the performance of an important token classifier in our setting. Here, we report the experiments that led us to use a linear classifier based on draft token embeddings encoded with both $\theta_{draft}$ and $\theta_{main}$. To that end, we compare the different classifier variants using the important token embeddings from the GSM8K [Cobbe et al., 2021] training subset (see Section 4.1).
@@
For the first set of experiments (Figure 6), we compare regularizer coefficients for Logistic Regression (left). We also report different classifier types: Logistic Regression, a Random Forest with 128 trees and a multi-layer perceptron (MLP) with a single hidden layer consisting of 128 hidden units with ReLU activation. For consistency, we run all models using scikit-learn [Pedregosa et al., 2011] v1.4.2 with all other settings kept to their default values. For LogisticRegression, we set max_iter parameter to $500$. For MLP, we perform early stopping on yet another 10% subset of the training set with the built-in default MLPClassifier early stopping parameters. For RandomForest, we use min_samples_leaf set to $0.001$. For this evaluation, all classifiers use draft and target model hidden states (concatenated) encoding the draft token, which is our main setup from Section 3.2.
@@ Appendix I On the Instability of vLLM Inference for Accuracy Benchmarks
In Section 4, we evaluate AutoJudge decoding in terms of the number of accepted tokens per phase and the real-world tokens per second. For convenience, we only use vLLM implementation to evaluate the number of tokens per second and report all other metrics based on transformers / pytorch implementation. This is because, in our preliminary experiments, we observed discrepancies between benchmark accuracy with vLLM inference and PyTorch. Upon further investigation, we found that vLLM inference sometimes produces inconsistent results. Namely, if we run vLLM greedy inference in bfloat16 precision with speculative decoding, changing “technical” hyperparameters such as window size can significantly affect the accuracy. To illustrate this, we measure GSM8K accuracy on 132 random test samples in the same setup as Section 4.3. We use standard vLLM implementation of speculative decoding Leviathan et al. [2023] without any modifications to the vLLM codebase.
@@
The results in Table  9 illustrate the observed inconsistency: changing window size can affect the accuracy to a significant degree. This coincides with concurrent observations about inconsistent LLM inference from Yuan et al. [2025] and  He and Lab [2025]. To make our measurements more consistent, we report all accuracy and accepted token rates using a more stable implementation based on Hugging Face Transformers (included in our repository), only using vLLM to measure the speed in terms of tokens/second.
@@
Following the invention of speculative decoding, several lines of work proposed follow-up algorithms that incorporate trained speculation “heads”[Cai et al., 2024, Li et al., 2024d, c], tree decoding [Miao et al., 2023, Chen et al., 2024, Svirschevski et al., 2024] and many other improvements. In this section, we explore how AutoJudge generalizes to these more advanced decoding algorithms.
To that end, we integrate AutoJudge with the popular EAGLE-2 algorithm [Li et al., 2024c]. Unlike the original speculative decoding, EAGLE-2 does not have a separate draft model, but trains a lightweight “head” to predict future tokens from target model hidden states. This allows EAGLE to draft tokens much faster, albeit less accurately than powerful standalone draft model.
@@
Case B: Creative Writing. We use the “Creative Writing” subset of Arena-Hard-Auto-v2.0 [Li et al., 2024b, a] that contains 250 creative writing tasks sourced from Chatbot Arena. These tasks include requests to write a poem on a certain topic, an imaginary dialogue transcript, or similar, in several languages. These tasks are even harder to judge than question answering, often boiling down to personal preference. We set 100 queries aside for evaluations and use the rest to mine important tokens, using the standard generation prompt from Arena-Hard-Auto-v2.
@@
When evaluating on creative writing, we found that the default LLM judge often produces different answers between consecutive runs, both with sampling and greedy decoding, likely due to numerical instability [Yuan et al., 2025]). To make our results more reliable we switch to a stronger claude-sonnet-4-20250514 judge model and run it 3 times with majority voting:
@@
To evaluate the task transfer capability of AutoJudge, we test whether a classifier trained on one mathematical reasoning dataset can generalize to a different, more challenging dataset. For this evaluation, we use the AutoJudge classifier trained on GSM8K 0-shot with the Llama-3.2-1B-Instruct draft / Llama-3.1-8B-Instruct target model pair (from Section 4.1) and apply it to the MATH-hard subset from the LLama Codebook ∗*∗**We reproduce the setup and get the problems from the official Meta evaluations repository. based on the lm-evaluation-harness benchmark [Gao et al., 2021]. The MATH-hard dataset contains significantly more challenging mathematical problems compared to GSM8K, requiring more advanced mathematical reasoning. We follow the standard evaluation protocol from the lm-evaluation-harness framework and report exact match accuracy as the downstream metric.
@@
Figure 15: GSM8K 0-shot evaluation of Llama-3.2-1B draft / Llama-3.1-8B target model pair comparing AutoJudge with manual annotation mining (following [Bachmann et al., 2025]) against Top-K baseline.
To compare our automated approach with the manual annotation procedure from Judge Decoding (Bachmann et al. [2025]), we conduct an experiment following their methodology as closely as possible. Unfortunately, the original paper does not provide public access to their annotated dataset or source code, which limits direct comparison. Therefore, we recreate their data collection and training procedure based on the methodology described in Section 4.1 of their paper.
@@
We curate a dataset of 167 high-quality questions from two sources: the ARC-Challenge benchmark Clark et al. [2018] (25 questions) and the Alpaca dataset Taori et al. [2023] (142 questions, split between 71 mathematical reasoning problems and 71 coding problems). Since the exact question indices used in the original Judge Decoding work are not publicly available, we manually selected challenging and diverse questions from each dataset that represent a range of difficulty levels similar to those described in the original paper.
@@
Following the procedure described in "Dataset Curation" in Section 4.1 of [Bachmann et al., 2025], we manually annotate each generated response to identify the precise location where errors occur. We encode all responses using the Llama-3.1-8B-Instruct model to extract hidden state representations, then manually label each token as either correct or incorrect based on whether accepting that token would lead to a wrong final answer. This annotation process required approximately 8-10 hours of careful manual review by the authors.
@@
To estimate the energy consumption of AutoJudge, vanilla speculative decoding, and sequential decoding, we run each inference method on a 10% sample of the GSM8K test set using vLLM. We measure real-world power usage on Llama 3.2 1B draft / 3.1 8B target models with a single A100-SXM4-80GB GPU (Watts) as reported by nvidia-smi, and multiply it by the mean inference time. This represents the GPU-reported power consumption, before adjusting for PSU inefficiency (equally for AutoJudge and baselines) and will vary between GPU types. For convenience, we convert all results to kJ, similar to Maliakel et al. [2025].
@@
Hence, we can run Algorithm 1 by replacing GENERATE(…) on lines 4 and 10 with a call to an LLM generation API with the specified input. With this reframing, we can mine important tokens by querying LLM API providers even if they cannot inference the large target model locally. This can help in a number of use cases: for instance, when developing a speculative decoding algorithm for use with offloading [Svirschevski et al., 2024, Miao et al., 2023].
@@
  • The GSM8K benchmark [Cobbe et al., 2021] is licensed under the MIT License;
  • The LiveCodeBench benchmark [Jain et al., 2024] is licensed under the MIT License;
  • Llama 3.1 and 3.2 models Dubey et al. [2024] are under the Llama Community License Agreement.
Our work would also be impossible without open-source software including (but not limited to) PyTorch [Paszke et al., 2019], Hugging Face Transformers [Wolf et al., 2019], vLLM [Kwon et al., 2023] and hundreds of other open-source packages in the Python data science & machine learning ecosystem. Enumerating and acknowledging all these individual packages would take up several pages, but they can be recovered automatically from our repository.

### v4

@@ 4.3 Inference Benchmarks with vLLM
We compare AutoJudge decoding against lossless vLLM speculative decoding (leviathan2023fast). To ensure fair comparison, we tune draft window size for AutoJudge and baseline independently (see Appendix G). In each configuration, we report the absolute speed in tokens per second (batch size 1), as well as the relative speedup over speculative decoding. We run 1B/8B model pair on a single A100-SXM4-80GB GPU; 8B/70B on 4 A100-SXM4-80G GPUs in tensor-parallel mode. Finally, the 8B/405B runs on 8 H100-SXM5-80GB GPUs with the 405B model loaded in FP8 precision. Additionally, we consider a setup where 8B/70B model pair runs on a single A100-SXM4-80GB GPU with RAM offloading (see Appendix H). We provide additional configuration details in Appendix E.
@@ 5 Conclusion
One promising direction for future research is to focus on open-ended problems such as instruction following or creative writing. While AutoJudge already demonstrates some speedups on open-ended tasks, it was primarily designed for technical problems such as math and programming. It would be interesting to investigate how it can be extended for longform generations with noisy and informal quality criteria. Another possible direction is to combine AutoJudge with additional speculative decoding algorithms, such as speculative decoding with tree-based drafts (specinfer; chen2024sequoiascalablerobusthardwareaware; specexec) or learned drafting heads (cai2024medusa; li2025eagle3scalinginferenceacceleration).
@@ Acknowledgements
In Section 3.1, we assume that the generation procedure is deterministic, i.e. that the model performs “greedy inference”. In practice, however, many applications work better with stochastic sampling [holtzman_nucleus_sampling]. However, this has an obvious caveat for Algorithm 1: if the text generation process is stochastic, a token can be deemed important based not on its actual impact on the model outputs, but on the randomness of the decoding procedure.
To generalize our approach for stochastic generation, we take advantage of the well-known Gumbel-max trick [gumbel1954statistical]. To recap, if we add independent Gumbel-distributed random variables to each predicted logit and take the index of the maximum, the probability that a certain index will be chosen is equal to the softmax of the original logits.
@@
As we described earlier, Algorithm 1 is inherently sequential because it searches not for individual important tokens, but for important token combinations. In principle, it is tempting to consider a simpler algorithm that considers each token replacement in isolation and can run in parallel. However, when considering [target_model_gen_0, draft_token, target_model_gen_1] sequences only, a sufficiently strong target model might recover from even a low-quality token and still produce the correct answer. This results in a failure mode where all tokens are individually unimportant, but when all such tokens are jointly replaced with their draft versions, the model fails to produce the correct answer. In our preliminary experiments, when using Llama-3.1-70B-Instruct touvron2023llama as the target model and Llama-3.2-1B-Instruct as the draft model, fewer than 1% of the tokens were labeled as important with this simplified algorithm, whereas our main Algorithm 1 found substantially more. One interesting guarantee of Algorithm 1 over its naïve counterpart is that whenever draft and target models produce different (non-equivalent) answers to a given prompt, our algorithm will find at least one important token, whereas the naïve algorithm may find none.
@@ Appendix B Additional Details on Classifier Training
As we discussed earlier in Section 3.2, there are several important design choices that can affect the performance of an important token classifier in our setting. Here, we report the experiments that led us to use a linear classifier based on draft token embeddings encoded with both $\theta_{draft}$ and $\theta_{main}$. To that end, we compare the different classifier variants using the important token embeddings from the GSM8K [cobbe2021gsm8k] training subset (see Section 4.1).
@@
For the first set of experiments (Figure 6), we compare regularizer coefficients for Logistic Regression (left). We also report different classifier types: Logistic Regression, a Random Forest with 128 trees and a multi-layer perceptron (MLP) with a single hidden layer consisting of 128 hidden units with ReLU activation. For consistency, we run all models using scikit-learn [pedregosa2011scikit] v1.4.2 with all other settings kept to their default values. For LogisticRegression, we set max_iter parameter to $500$. For MLP, we perform early stopping on yet another 10% subset of the training set with the built-in default MLPClassifier early stopping parameters. For RandomForest, we use min_samples_leaf set to $0.001$. For this evaluation, all classifiers use draft and target model hidden states (concatenated) encoding the draft token, which is our main setup from Section 3.2.
@@ Appendix I On the Instability of vLLM Inference for Accuracy Benchmarks
In Section 4, we evaluate AutoJudge decoding in terms of the number of accepted tokens per phase and the real-world tokens per second. For convenience, we only use vLLM implementation to evaluate the number of tokens per second and report all other metrics based on transformers / pytorch implementation. This is because, in our preliminary experiments, we observed discrepancies between benchmark accuracy with vLLM inference and PyTorch. Upon further investigation, we found that vLLM inference sometimes produces inconsistent results. Namely, if we run vLLM greedy inference in bfloat16 precision with speculative decoding, changing “technical” hyperparameters such as window size can significantly affect the accuracy. To illustrate this, we measure GSM8K accuracy on 132 random test samples in the same setup as Section 4.3. We use standard vLLM implementation of speculative decoding leviathan2023fast without any modifications to the vLLM codebase.
@@
The results in Table  9 illustrate the observed inconsistency: changing window size can affect the accuracy to a significant degree. This coincides with concurrent observations about inconsistent LLM inference from yuan2025fp32deathchallengessolutions and  he2025nondeterminism. To make our measurements more consistent, we report all accuracy and accepted token rates using a more stable implementation based on Hugging Face Transformers (included in our repository), only using vLLM to measure the speed in terms of tokens/second.
@@
Following the invention of speculative decoding, several lines of work proposed follow-up algorithms that incorporate trained speculation “heads”[cai2024medusa, li2024eagle, eagle2], tree decoding [specinfer, chen2024sequoiascalablerobusthardwareaware, specexec] and many other improvements. In this section, we explore how AutoJudge generalizes to these more advanced decoding algorithms.
To that end, we integrate AutoJudge with the popular EAGLE-2 algorithm [eagle2]. Unlike the original speculative decoding, EAGLE-2 does not have a separate draft model, but trains a lightweight “head” to predict future tokens from target model hidden states. This allows EAGLE to draft tokens much faster, albeit less accurately than powerful standalone draft model.
@@
Case B: Creative Writing. We use the “Creative Writing” subset of Arena-Hard-Auto-v2.0 [arenahard2024, li2024crowdsourced] that contains 250 creative writing tasks sourced from Chatbot Arena. These tasks include requests to write a poem on a certain topic, an imaginary dialogue transcript, or similar, in several languages. These tasks are even harder to judge than question answering, often boiling down to personal preference. We set 100 queries aside for evaluations and use the rest to mine important tokens, using the standard generation prompt from Arena-Hard-Auto-v2.
@@
When evaluating on creative writing, we found that the default LLM judge often produces different answers between consecutive runs, both with sampling and greedy decoding, likely due to numerical instability [yuan2025fp32deathchallengessolutions]). To make our results more reliable we switch to a stronger claude-sonnet-4-20250514 judge model and run it 3 times with majority voting:
@@
To evaluate the task transfer capability of AutoJudge, we test whether a classifier trained on one mathematical reasoning dataset can generalize to a different, more challenging dataset. For this evaluation, we use the AutoJudge classifier trained on GSM8K 0-shot with the Llama-3.2-1B-Instruct draft / Llama-3.1-8B-Instruct target model pair (from Section 4.1) and apply it to the MATH-hard subset from the LLama Codebook ∗*∗**We reproduce the setup and get the problems from the official Meta evaluations repository. based on the lm-evaluation-harness benchmark [eval-harness]. The MATH-hard dataset contains significantly more challenging mathematical problems compared to GSM8K, requiring more advanced mathematical reasoning. We follow the standard evaluation protocol from the lm-evaluation-harness framework and report exact match accuracy as the downstream metric.
@@
Figure 15: GSM8K 0-shot evaluation of Llama-3.2-1B draft / Llama-3.1-8B target model pair comparing AutoJudge with manual annotation mining (following [bachmann2025judgedecoding]) against Top-K baseline.
To compare our automated approach with the manual annotation procedure from Judge Decoding (bachmann2025judgedecoding), we conduct an experiment following their methodology as closely as possible. Unfortunately, the original paper does not provide public access to their annotated dataset or source code, which limits direct comparison. Therefore, we recreate their data collection and training procedure based on the methodology described in Section 4.1 of their paper.
@@
We curate a dataset of 167 high-quality questions from two sources: the ARC-Challenge benchmark arc_allenai (25 questions) and the Alpaca dataset alpaca (142 questions, split between 71 mathematical reasoning problems and 71 coding problems). Since the exact question indices used in the original Judge Decoding work are not publicly available, we manually selected challenging and diverse questions from each dataset that represent a range of difficulty levels similar to those described in the original paper.
@@
Following the procedure described in "Dataset Curation" in Section 4.1 of [bachmann2025judgedecoding], we manually annotate each generated response to identify the precise location where errors occur. We encode all responses using the Llama-3.1-8B-Instruct model to extract hidden state representations, then manually label each token as either correct or incorrect based on whether accepting that token would lead to a wrong final answer. This annotation process required approximately 8-10 hours of careful manual review by the authors.
@@
To estimate the energy consumption of AutoJudge, vanilla speculative decoding, and sequential decoding, we run each inference method on a 10% sample of the GSM8K test set using vLLM. We measure real-world power usage on Llama 3.2 1B draft / 3.1 8B target models with a single A100-SXM4-80GB GPU (Watts) as reported by nvidia-smi, and multiply it by the mean inference time. This represents the GPU-reported power consumption, before adjusting for PSU inefficiency (equally for AutoJudge and baselines) and will vary between GPU types. For convenience, we convert all results to kJ, similar to maliakel2025investigatingenergyefficiencyperformance.
@@
Hence, we can run Algorithm 1 by replacing GENERATE(…) on lines 4 and 10 with a call to an LLM generation API with the specified input. With this reframing, we can mine important tokens by querying LLM API providers even if they cannot inference the large target model locally. This can help in a number of use cases: for instance, when developing a speculative decoding algorithm for use with offloading [specexec, specinfer].
@@
  • The GSM8K benchmark [cobbe2021gsm8k] is licensed under the MIT License;
  • The LiveCodeBench benchmark [jain2024livecodebenchholisticcontaminationfree] is licensed under the MIT License;
  • Llama 3.1 and 3.2 models dubey2024llama are under the Llama Community License Agreement.
Our work would also be impossible without open-source software including (but not limited to) PyTorch [pytorch], Hugging Face Transformers [wolf2019huggingface], vLLM [kwon2023efficient] and hundreds of other open-source packages in the Python data science & machine learning ecosystem. Enumerating and acknowledging all these individual packages would take up several pages, but they can be recovered automatically from our repository.